The length l of each side of a square is decreased by 2 and a half units. The perimeter of the square with the new side lengths is 66 units.
What was the original side length?

Answer by ewatrrr(24785)


Hi
P = 4(L + 2.5) = 66 Perimeter is the Distance ALL the way around the new Square
4L + 10 = 66
4L = 56
L = 14 units, the original side length